OpenStudy (anonymous):

Perform the operation and write the answer in simplest form 6/10 / 12/7

OpenStudy (texaschic101):

dividing fractions... take the number you are dividing by, and " flip " it, then multiply. (6/10) / (12/7) = 6/10 * 7/12 = 42/120 which reduces to 7/20

OpenStudy (anonymous):

This is exactly what I did, and the test still said I was wrong.

OpenStudy (amilapsn):

\[6\div10\div12\div7=\frac{6}{10\times12\times7}\]

OpenStudy (amilapsn):

It can be thought in this way too :)

OpenStudy (anonymous):

Well thanks guys.